OpenStudy (anonymous):

If the perimeter of a rectangular house is 44 yards, and the length is 36 feet, what is the width of the house? A. 10 yards B. 18 yards C. 28 feet D. 32 feet E. 36 yards

OpenStudy (anonymous):

.If the perimeter of a rectangular house is 44 yards and the length is 36 feet. What is the width of the house? We can either change yards to feet, or feet to yards...<< 1. Changing yards to feet. Perimeter = 44 yards = 44(3 feet) = 132 feet Perimeter = 2(Length + Width) P = 2(L + W) 132 = 2(36 + W) 132 = 72 + 2W 132 - 72 = 2W 60 = 2W 30 = W So the width is 30 feet.

OpenStudy (anonymous):

Changing feet to yards. Length = 36 feet = 36(1/3 yard) = 12 yards Perimeter = 2(Length + Width) P = 2(L + W) 44 = 2(12 + W) 44 = 24 + 2W 44 - 24 = 2W 20 = 2W 10 = W So the width is 10 yards.

OpenStudy (anonymous):

kumar2006 y didnt u help me mine was easier u just pooped on me